19. Differential Growth Synovec Corp. is experiencing rapid growth. Dividends are expected to grow at 27 percent per year during the next three years, 18 percent over the following year, and then 4 percent per year indefinitely. The required return on this stock is 10 percent, and the stock currently sells for $71 per share. What is the projected dividend for the coming year?

I am trying to create an MS Excel table to help me with the various calculations for this question.
Answer #1
The price of the share of $71 is the PV of the
expected dividends discounted at the required
rate of return of 10%.
Let the last dividend be 'a'.
Then the sum of the PV of the 1st 3 years' dividend = a*1.27/1.1+a*1.27^2/1.1^2+a*1.27^3/1.1^3 = a*(1.27/1.1+1.27^2/1.1^2+1.27^3/1.1^3) =
= a*4.0265
PV of the 4th year's dividend = a*(1.27^3*1.18)/1.1^4 = a*1.6509
PV of the dividends form 5th year = a*1.27^3*1.18*1.04/((0.1-0.04)*1.1^4) = a*28.6157
Therefore, $71 = a*(4.0265+1.6509+28.6157)
71 = a*34.2931
a = 71/34.2931 =
a [last year dividend] = $2.07
Projected dividend for the coming year = 2.07*1.27 = $                 2.63